If Phumza and Nathi eat together, they can eat 480 hamburgers in 6 hours. However, eating alone, Nathi takes 16 hours longer than Phumza would eat 480 hamburgers.
How long does it take Phumza to eat 480 hamburgers?
1
Expert's answer
2021-10-04T17:08:08-0400

Let's say phumba can eat ""x"" hamburger while Nathi can eat ""y"" hamburger

Time taken 6 hours

Therefore we can say together they can eat

"6(X+y)=480....(1)"

Phumba alone can eat for ""t"" hours

"tx=480\\t=\frac{480}{t}"

Nathi alone can eat

"(t+16)y=480\\t=\frac{480}{t+16}"

Substitute values of X and y in equition 1

"6(\frac{480}{t}+\frac{480}{t+16}=480\\\frac{480}{t}+\frac{480}{t+16}=80\\\frac{480(t+16)+480t}{t(t+16)}=80"

"=480(t+16)+480t=80t(t+16)\\=6(t+16)+6t=y(t+16)\\=6t+96+6t=t^2+16t\\0=t^2+16t-12t-96\\0=t^2+4t-96\\t^2+4t-96=0"


"t^2+12t-8t-96 =0\\t(t+12)-8(t+12)=0\\(t+12)(t-8)=0\\t=-12,t=8"


t cannot be negative, therefore t=8

Phumba takes 8 hours to eat 480hamburgers